• USGS
  • NSF
  • Iris
  • Iris

Recent Maintenance - NNA

NNA operational status: 

2020.07.02

The station is back online as of day 2020,184-16:59 after the station operator replaced the Newmar power supply and backup batteries.

 

2020.05.22

The station is down as of day 2020,139-19:47 due to failure of the Newmar power supply. 

 

2020.01.31

The station was down from day 2020,028-17:34 to 2020-031-20:12 due to power problems (Newmar power supply and backup battery failures).  The station operator replaced the backup batteries, however output from the power supply is still low.  IDA engineers are working with the station operator to resolve this issue.

 

2018.08.17

The STS-1 channels were restored as of 2018,229-18:43 (Aug. 17) after the station operator replaced the E300.

2018.03.28

The STS-1 channels are non-seismic as of day 2018,087 after the E300 was disconnected and sent for repair.

2017.01.20

Communication between the station computers and digitizers was restored after an IDA engineer visited the station and replaced the ethernet switch on 2017,016.

2015.11.10

The station is back up as of day 2015,314-1829 after the station operator visited the station and replaced the surge arrestor and circuit breaker with equipment shipped from IDA/UCSD.

 

2015.10.08

The station is down as of day 2015,274-1720 due to damage to the power system caused by a power surge (presumably from lightning).  IDA engineers are working with the station operator to resolve this issue.

 

2015.08.24

The station is back online as of day 236-1936 (Aug.24) after an IDA engineer visited the station and replaced the station computers, and upgraded the power system. A Setra microbarometer was also installed.

 

2015.08.20

The station is still down. An IDA engineer will be onsite next week to resolve this issue and upgrade the power system.

 

2015.08.13

The station operator reported the station is down (as of day 218-1207) due to a power supply failure.  The station operator is planning on visiting the station to replace the power supply.

 

2011.03.03

The station has been upgraded to a Quanterra Q330HR system. Sensors include an STS-1, STS-2, and Episensor strong motion sensor.

2010.11.08

The station was down from day 2010,304-06:28 to 2010,308-09:45 due to problems with the DAS. Data acquisition was restored after the station operator cycled power to the DAS.

2010.10.06

The station is back on-line as of day 2010,278. However, the barometer channel (wdi00) is flat-lined. A station upgrade is being planned.

2010.09.22

The station is down as of day 2010,263-12:54 due to problems with the DAS. The station operator has been requested to cycle power to the DAS.

2010.08.10

The station was down from day 2010,221-11:17 to 2010,222-21:08 due to problems with the DAS.

2010.06.01

The station is back on-line as of day 2010,151-16:56 after the station operator cycled power to the DAS.

2010.05.28

IDA station NNA is down since day 2010,147-15:50 due to problems with the DAS. The station operator has been requested to visit the station to investigate, and cycle power to the DAS.

2010.03.29

The station was down from day 2010,084-23:51 to 2010,088-00:10 due to problems with the DAS. The station came back online after repeated attempts to cycle power to the DAS.

2009.12.01

The station is back up as of day 2009,335-21:36 after the station operator reconnected a cable from the digitizer to a switch onsite.

2009.11.27

The station is down as of day 2009,331-01:14 due to problems with the digitizer that are being investigated.

2009.11.03

The station was down from day 2009,305-05:35 to 2009,307-12:51 due to a power outage. Data acquisition resumed once power was restored and the station computer rebooted.

2009.10.30

The station is back online as of day 2009,303 after the station operator replaced the digitizer.

2009.09.30

The station is down again since day 2009,273-14:20 due to problems with the digitizer.

2009.09.28

The station is back on-line as of day 2009,271. However, problems with the digitizer are intermittent and a new digitizer is being shipped to the station.

2009.09.24

The station is down as of day 2009,265-04:22 due to problems with the digitizer. A new digitizer is being prepared to ship to the station.

2009.08.31

The station is currently up. However, problems persist with the UPS and digitizer following the site visit by an IDA engineer. Plans for a replacement digitizer are being prepared.

2009.08.19

The station is back on-line as of day 2009,231 after an IDA engineer visited the station and replaced the power supply and CPU stack for the DAS. The engineer is currently onsite troubleshooting problems with the DAS.

2009.08.05

The station is down as of day 2009,217 due to problems with the power system and the DAS.

2009.06.15

All channels from the STS-2 seismometer are non-seismic since about day 2009,167. Testing is currently underway to investigate excessive spiking on all horizontal channels at the station.

2009.04.18

The horizontal channels from both the STS-1 and STS-2 seismometers are spiking since about day 2009,108 for reasons yet to be determined.

2007.11.05

The station is back online after an IDA engineer visited the station and reconfigured the GPS clock.

2007.05.01

The station is down due to problems with the GPS clock and the DAS that are yet to be determined.

2007.01.18

The GPS clock went back into a locked state and normal operation resumed as of day 018-23:37, with no loss of data.

2006.10.21

The GPS clock failed around day 294-11:48. A new clock is being prepared to send to the station.

2006.06.13

The data acquisition system has been upgraded to a MK8 system. New equipment including an STS-2 seismometer, strong motion episensor, ISP, GPS clock, power supplies, radios, and a Paroscientific microbarograph have been installed.

2006.04.07

The station is back up after the station operator repaired the power supply.

2006.03.27

The station is down due to failure of the remote power supply. A station visit is planned to restore operation and upgrade the station.

2005.04.11

The station is up as of 4/09 after an IDA engineer visit to the site to repair the station. The problem was corrected by replacing a DC/DC converter for the DAS power supply.

2005.03.28

The station is down as of 3/21 due to problems with the DAS.

2005.02.18

An IDA engineer visited the site and replaced the GPS clock and DAT drive (02/17/05).

2005.01.05

Station operator reports GPS clock failure (01/03/05). Station was down from day 12/08/04 to 12/13/04 due to problems with the DAS. Data acquisition resumed after station operator replaced circuit boards in the DAS and ARS.

2004.06.01

Telemetry data are avilable after the telemetry relay housing has been rebuilt.

2004.04.23

Operator reports radios used to telemeter data have been vandalized and will require at least one week to repair.

2004.02.17

Telemetry restored after station operator installed new radios.

2003.09.23

Station operable again after tape drive and power supply board were fixed on August 8th.

2003.01.22

Operator reports telemetry relay station has been vandalized. Telemetry to be down indefinitely.

2002.07.22

New EPROMs installed to correct DSP problem.

2002.04.22

Station operating normally.

On 2002.02.28, the operator repaired the telemetry link to the station using spare parts on site. Thieves had stolen the spread spectrum radio and buffering hardware from station five days earlier.